Suppose that the economy is in recession with a recessionary gap of $1 trillion. The MPC is 0.9 and the tax rate on income is 30%. Answer the following
(a) Compute the value of the multiplier.
(b) Suppose the AS curve is completely flat. What would be the change in G needed to make the gap equal to zero?
(c) Would the change in G you found in b. be enough if the AS curve was positively sloped? Explain using the appropriate graph.
1
Expert's answer
2016-05-16T10:47:02-0400
Recessionary gap of $1 trillion, c = 0.9, t = 0.3.
(a) The value of the multiplier is m = 1/(1 - c*(1 - t)) = 1/(1 - 0.9*(1 - 0.3)) = 1/0.37 = 2.7.
(b) If the AS curve is completely flat, then the change in G needed to make the gap equal to zero is changeY/changeG = m, so:
1/changeG = 2.7, change in G = 1/2.7 = $0.37 trillion.
(c) If the AS curve was positively sloped, then this change may be lower than it is needed.